OpenStudy (anonymous):

megan bought 7 charms for 31.50$. Each charms costs the same amount of money. Write an inequality that can be used to find the maximum amount of charms(c) Megan can buy with $75 ________ what is the maximum amount of charms Megan can buy with $75

OpenStudy (anonymous):

they are two seperate questions thats what the line is for

OpenStudy (anonymous):

no ria please help meeeee

OpenStudy (ria23):

31.50/7 = 4.5 each 4.5x ≤ 75

OpenStudy (anonymous):

but what formula do i use

OpenStudy (ria23):

Are these what yhu are looking for? I could explain further if needed. c:

OpenStudy (ria23):

ok, one sec.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

like c=222a+44 something like that idk

OpenStudy (ria23):

31.50/7 = x 75/x= y X= individual charm cost Y= how many charms she can buy

OpenStudy (ria23):

Does this help.

OpenStudy (anonymous):

ok uhm not really

OpenStudy (anonymous):

i have to use c for a vairable

OpenStudy (ria23):

Alright. $31.50 is the amount of money she used to buy 7 charms.

OpenStudy (anonymous):

yes

OpenStudy (ria23):

Yhu want to take the amount She used to buy the charms, and divide that by the amount of charms she bought. giving yhu the equation 31.50/7= x x being the amount she spent on the charms. Doing that, what do yhu get.

OpenStudy (anonymous):

ok what is the second part of the answer

OpenStudy (anonymous):

also -2x-1=3x+5 solve for x ---------- 2

OpenStudy (ria23):

Ok, so the next step, yhu get 4.5 for x. So it'll be 4.5x ≤ 75 or 4.5c ≤ 75

OpenStudy (ria23):

Give me just a second with the second question.

OpenStudy (anonymous):

ok

OpenStudy (ria23):

Yhu have the original questions answered, correct?

OpenStudy (anonymous):

yes this is a new question number 7 that was 6

OpenStudy (ria23):

ok, Is there multiple choice options, or is it a work out the problem and show yhur work kind of problem.

OpenStudy (anonymous):

show work and write out answer

OpenStudy (ria23):

 First, subtract the 3x from both sides.